Question 4 20 pts You plan to give your child a new car for her graduation in 18 months. The car costs $23,000 How much mustplease answer
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

r = 2.45%/12 = 0.2042% (monthly)

$23,000 = PV * (1 + 0.002042)18

$23,000 = PV * 1.0374

PV = $22,170.92